The Landers Corporation needs to raise $1.60 million of debt on a 20-year issue. If it places the bonds privately, the interest rate will be 10 percent. Twenty thousand dollars in out-of-pocket costs will be incurred. For a public issue, the interest rate will be 9 percent, and the underwriting spread will be 2 percent. There will be $120,000 in out-of-pocket costs. Assume interest on the debt is paid semiannually, and the debt will be outstanding for the full 20-year period, at which time it will be repaid. Use Appendix B and Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods.


a. For each plan, compare the net amount of funds initially available—inflow—to the present value of future payments of interest and principal to determine net present value. Assume the stated discount rate is 12 percent annually. Use 6 percent semiannually throughout the analysis. (Disregard taxes.) (Assume the $1.60 million needed includes the underwriting costs. Input your present value of future payments answers as negative values. Do not round intermediate calculations and round your answers to 2 decimal places.)

Private Placement    Public Issue   
Net Amount to Landers:
Present Value of Future Payments:
Net Present Value:

  

b. Which plan offers the higher net present value?

  • Private placement

  • Public issue
Answer #1

А в 1 a) 2 Net amount to Landers Private placement 1,580,000 =1600000-20000 $ Public Issue 1,448,000 =1600000-32000-120000 Pr

.Please rate thumbs up